OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? the line perpendicular to y = 1/3x + 5 through (2, 1) A. y = -1/3x + 7 B. y = 3x + 7 C. y = 1/3x + 7 D. y = -3x + 7

OpenStudy (anonymous):

i know its not A

OpenStudy (anonymous):

Line perpendicular to y = 1/3x + 5 m = 1%2F3 through (2, 1) m = -3 y = -3x + b 1 = -3 * 2 + b Using (x, y) ordered pair (2, 1) to solve for b. 7 = b y = -3x + 7

OpenStudy (anonymous):

THANK YOU!!!

OpenStudy (anonymous):

You're welcome!! Xo!